(5+√6)^x-2 +(5-√6)^x-2=10Finf the value of x Plz reply fast i need the solution of tbis questine

Grade:10

2 Answers

Nishant Vora IIT Patna
7 years ago
5+√6 = 19 /5 – √6

hence

(5+√6)^x-2 +(5-√6)^x-2=10
=> (19 /5 –√6 )^ x-2 +(5-√6)^x-2=10
let(5-√6)^x-2 = t and you will get a quadratic eqn now solve
Amoolya
7 years ago
Sir what about 19^x-2Pls send the full solution......The equation left ia 19^x-2=10t-t^2What should i do next to solve this ques